# Dichotomous Diffusion Policy Optimization

### v1

@@ Abstract
Project Page: https://lrmbbj.github.io/DIPOLE/
@@ 1 Introduction
Due to the strong capability of diffusion models (sohldickstein2015deep; ho2020denoising) in modeling multi-modal action distributions and controllable generation during inference (dhariwal2021diffusion; ho2022cfg), modeling policies using diffusion models has become a popular choice in solving complex decision-making tasks such as embodied robotics (chi2023diffusion; zheng2025x) and autonomous driving (zheng2025diffusionplanner; tan2025flow; li2025discrete). Although proven to be effective in imitation learning-based settings, training large diffusion/flow matching policies that surpass data-level performance with reinforcement learning (RL) (sutton1998reinforcement) has remained an important yet challenging direction.
@@
Training diffusion policies with RL faces numerous challenges, most notably, learning stability and computation efficiency. A naïve approach to train diffusion policies with RL is to directly optimize the reward or value objective via gradient backpropagation through the multi-step denoising process (xu2023imagereward; clark2023directly), which often suffers from noisy and unstable gradient updates, while also being extremely costly. To avoid this, some studies adopt a compromise by freezing the diffusion model and instead searching for optimized noises (wagenmaker2025steering; hansen2023idql), a strategy often referred to as inference-time scaling (ma2025inference). However, these approaches rely heavily on well-pretrained diffusion policies and are fundamentally limited by their performance upper bound. Another explored direction is to adopt policy gradient methods (such as PPO (schulman2017proximal)) for diffusion policy optimization, which models the denoising process as a multi-step Markov decision process (MDP) and uses Gaussian approximations to compute the log-likelihood of intermediate denoising steps (black2023training; ren2024diffusion). However, the crude Gaussian-based approximation only provides reasonable likelihood information when adopting sufficiently small denoising steps, which inevitably results in large exploration spaces and prolonged training, making such methods difficult to scale and prone to approximation error accumulation in practice. Therefore, a critical research question arises: Can we build a more effective and stable RL method for diffusion policy optimization?
@@
To answer this question, we turn our attention to the KL-regularized RL objective, which offers a nice, closed-form weighted regression objective for optimal policy extraction (peng2019advantage). We can thus optimize a diffusion policy by incorporating an exponential reward- or value-based weighting term, scaled by a temperature parameter, into the standard diffusion regression loss (lee2023aligning; kang2023efficient; zheng2024safe). Although promising, this approach also suffers from several limitations. A fundamental issue is that weighted regression can only achieve greedy reward maximization when the temperature parameter is set to a large value, which easily leads to exploding loss and training instability. Moreover, the learning loss becomes dominated by a small number of high-reward samples, which severely undermines training effectiveness and scalability even with increased data (park2024value). To address the previous challenges, we propose DIPOLE (Dichotomous diffusion Policy improvement), a novel RL framework designed for highly stable and controllable diffusion policy optimization. Specifically, we introduce a greedified KL-regularized RL objective, which regularizes policy learning towards a value-reweighted reference policy. Interestingly, we show that the original unstable exponential weighting term in the optimal policy can be decomposed into two bounded smooth dichotomous terms. This naturally allows us to decompose the optimal policy into a pair of stably learned dichotomous policies: one aims at reward maximization and the other focuses on reward minimization. Moreover, the optimized policy can be recovered through a linear combination of the scores from both dichotomous policies, which closely aligns with the widely used classifier-free guidance mechanism in diffusion models (ho2022cfg), enabling perfect controllability over the greediness of action generation.
@@
Extensive experimental results demonstrate the effectiveness of DIPOLE across a wide range of locomotion and manipulation tasks in ExORL (yarats2022exorl) and OGBench (ogbench_park2025) benchmarks, evaluated under both offline and offline-to-online RL settings. Furthermore, we scale our learning approach to a large vision-language-action (VLA) model and evaluate it on the large-scale real-world autonomous driving benchmark NAVSIM (Dauner2024NEURIPS), showcasing significant performance improvements over the pre-trained baseline. These results highlight the strong applicability of DIPOLE for complex, real-world decision-making scenarios.
@@ 2 Preliminary
Diffusion/flow matching policies. Diffusion and flow matching models have attracted significant attention due to their strong expressiveness in capturing multi-modal data distributions, making them popular policy classes for complex decision-making tasks such as robotics (chi2023diffusion; black2024pi0visionlanguageactionflowmodel) and autonomous driving (zheng2025diffusionplanner; tan2025flow). The action generation can be formulated as a state-conditional generation problem in which a probability path transforms a source distribution (typically a standard Gaussian) into a target action distribution. A neural network $\epsilon_{\theta}$ is trained to predict the noise along the path using the objective over a given dataset $\mathcal{D}$:
@@
where $a_{t}=\alpha_{t}a+\sigma_{t}\epsilon$ (we use the subscript $t$ to distinguish diffusion steps from MDP steps $k$), with $\alpha_{t}$ and $\sigma_{t}$ being predefined noise schedules commonly used in score-based diffusion models (song2021scorebased) or flow matching models (lipman2022flow). The multi-step diffusion process endows diffusion models with strong distribution-fitting capabilities. However, it also poses challenges for RL fine-tuning: gradient propagation through the entire diffusion process is costly and unstable; the exact likelihood computation with diffusion models is intractable, causing a series of problems when optimizing with existing policy gradient RL methods due to approximation error.
@@ 3.1 KL-Regularized Objective in RL
where $\beta>0$ is the temperature parameter, and $D_{\text{KL}}\left(p\|q\right)=\mathbb{E}_{x\sim p}\left[\log\left(p(x)/q(x)\right)\right]$. $G(\cdot)$ is the evaluated return to be maximized, which can either be the reward function $r(s,a)$ as in single-step problems such as LLM RL fine-tuning (korbak2022rl; shao2024deepseekmath), or the action-value function $Q^{\pi}(s,a)$ or advantage function $A^{\pi}(s,a)$ as in standard multi-step settings. The specific choice of reference policy $\mu$ gives rise to different RL task settings. For example, setting $\mu$ to be the uniform distribution, we recover maximum entropy RL as in SAC (haarnoja2018soft); setting $\mu$ to be the behavior policy in offline datasets $\mathcal{D}$, we obtain many offline RL algorithms (wu2019behavior; xu2023offline; gargextreme); lastly, setting $\mu$ to be a pre-trained policy $\pi_{0}$ or the recently updated policy $\pi_{k-1}$, it corresponds to offline-to-online fine-tuning scenarios (nakamoto2023cal; li2023proto) or trust-region style online policy optimization (schulman2015trust).
The flexibility of the KL-regularized RL framework makes it an ideal choice for diffusion policy optimization. The best part is, it is known that the optimization objective in Eq. (2) also provides a closed-form solution for optimal policy $\pi^{\star}$ as follows (nair2020awac):
@@
Intuitively, the optimal policy is a reweighted version of the reference policy $\mu$, in which actions with higher values are assigned greater probability density. As shown in several existing studies  (kang2023efficient; zheng2024safe), if given a pre-trained diffusion policy $\epsilon_{\theta}$ trained with Eq. (1) as the reference policy $\mu$, we can further optimize it with the weighted diffusion loss in Lemma 1 to extract the optimal diffusion policy $\epsilon^{\star}$.
@@
We can generate optimal $a\sim\pi^{\star}(a|s)$ in Eq. (3) by optimizing the weighted diffusion loss in Eq. (4) and solving the diffusion reverse process with obtained $\epsilon^{\star}$ (zheng2024safe).
@@
Compared to diffusion-based RL methods that rely on unstable reward/value maximization (xu2023imagereward; clark2023directly) or biased likelihood approximation (black2023training; ren2024diffusion), Eq. (4) offers a simple and scalable training scheme for policy optimization, requiring only the addition of a weighted term to the base diffusion learning objective in Eq. (1). Despite its simplicity, we do not observe the adoption of this scheme in many recent diffusion-based RL methods. Why is that? Actually, there exist several limitations for this exp-weighted regression scheme:
  • Optimality-stability trade-off. As the exponential function $\exp(\cdot)$ grows rapidly, a high-quality action with a large $G(s,a)$ value can lead to an extremely large weight term when $\beta$ is large, causing the explosion of learning loss and destabilizing the training process (illustrated in Figure 1). In practice, many methods mitigate this issue by either using a small $\beta$ or clipping the weighting term (gargextreme; xu2023offline; hansen2023idql). However, these treatments compromise the optimality of the extracted policy.
  • Inefficient learning. The training loss becomes dominated by a small number of high-return samples, which is inefficient for policy optimization (park2024value). Additionally, poor-quality samples still retain positive weight, which can adversely affect policy learning. The constrained optimization objective also makes the learning process highly dependent on the quality of the reference policy $\mu$, thereby limiting the potential for greedy policy optimization.
@@ 3.2 Dichotomous Diffusion Policy Improvement
In this revised objective, we instead regularize policy $\pi$ with a greedified, value-aware reference policy weighted by $\color[rgb]{0.76953125,0.1953125,0.19921875}\sigma\left(\beta G(s,a)\right)/Z(s)$}, where $Z(s)$ denotes the normalization factor and $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. This design shares a similar spirit with some offline RL methods that enhance policy performance by regularizing towards a greedier behavior policy or reward-weighted datasets (singh2022offline; hong2023beyond; xu2025optimal). It is worth noting that we use a bounded and smooth sigmoid function as the weighting function, which greedily assigns high weights to high-return samples while avoiding numerical instability. Moreover, we introduce a new hyperparameter $\color[rgb]{0.76953125,0.1953125,0.19921875}\omega$}, termed the greediness factor, which provides an additional interface for adjusting the greediness of policy extraction. We will reveal its role in the later derivation. Based on the optimization objective in Eq. (5), we can get its closed-form solution as follows:
@@
where $C$ is a constant. This shows that the score function of the optimal policy $\pi^{\star}$ can be expressed as a linear combination of scores of the dichotomous policies, weighted by $\omega$. Due to the inherent connection between the score function and the noise predictor in diffusion model (ho2020denoising), we can use $\tilde{\epsilon}\left(a_{t},s,t\right)=(1+w)\epsilon^{+}_{\theta_{1}}\left(a_{t},s,t\right)-w\epsilon_{\theta_{2}}^{-}\left(a_{t},s,t\right)$ in the reverse process of diffusion or flow matching for action sampling.
Interestingly, the formulation in Eq. (3.2) is remarkably similar to classifier-free guidance (CFG) (ho2022cfg), a popular method for enhanced conditional diffusion generation, which has the form of $\tilde{\epsilon}(x_{t},c,t)=(1+\omega)\epsilon_{\theta}(x_{t},c,t)-\omega\epsilon_{\theta}(x_{t},t)$, where $\epsilon_{\theta}(x_{t},c,t)$ is a conditioned version of $\epsilon_{\theta}(x_{t},t)$ with conditioning signal $c$. This reveals the inherent connection between our greedified KL-regularized RL objective and the CFG mechanism. Intuitively, our method further strengthens the positive distribution by pushing the negative distribution in the opposite direction, thus enabling flexible control of the optimality level of generated actions with the greediness factor $\color[rgb]{0.76953125,0.1953125,0.19921875}\omega$} (see illustration in Figure 1). Our final formulation also has some similarity with CFGRL (frans2025diffusion), which can be perceived as setting $\pi^{+}\propto\mu\cdot\mathbb{I}_{A\geq 0}$ and $\pi^{-}=\mu$ (AA is the advantage function). However, their method lacks theoretical backing, and using identical weights for both positive and negative samples limits the greediness of policy optimization, leading to suboptimal performance.
@@ 3.3 Practical Implementations
End-to-end autonomous driving. We also implement DIPOLE to train a large end-to-end autonomous driving model to demonstrate its scalability to solve real-world complex tasks. Specifically, we employ a non-reactive pseudo-closed-loop simulation based on real-world datasets for policy training. In this setup, the return $G(s,a)$ is defined by a reward function that evaluates trajectory quality based on safety, progress, and comfort. We employ a vision-language model (Florence-2 (xiao2024florence)) as the encoder and a diffusion action head as the decoder (zheng2025diffusionplanner). The model processes images from the left-front, front, and right-front cameras, along with language instructions such as ”turn left”, ”turn right”, and ”go straight”. This architecture results in a 1-billion parameter model, which we name DP-VLA, and is pre-trained using imitation learning. Subsequently, two separate LoRA modules are applied to the decoder to construct the positive and negative policies, allowing us to leverage Eq. (9) for training. We follow the offline-to-online RL setting to fine-tune the VLA model. Further implementation details are provided in Appendix E.

### v2

@@ 1 Introduction
Due to the strong capability of diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) in modeling multi-modal action distributions and controllable generation during inference (Dhariwal and Nichol, 2021; Ho and Salimans, 2022), modeling policies using diffusion models has become a popular choice in solving complex decision-making tasks such as embodied robotics (Chi et al., 2023; Zheng et al., 2025a) and autonomous driving (Zheng et al., 2025c; Tan et al., 2025; Li et al., 2025). Although proven to be effective in imitation learning-based settings, training large diffusion/flow matching policies that surpass data-level performance with reinforcement learning (RL) (Sutton et al., 1998) has remained an important yet challenging direction.
@@
Training diffusion policies with RL faces numerous challenges, most notably, learning stability and computation efficiency. A naïve approach to train diffusion policies with RL is to directly optimize the reward or value objective via gradient backpropagation through the multi-step denoising process (Xu et al., 2023b; Clark et al., 2023), which often suffers from noisy and unstable gradient updates, while also being extremely costly. To avoid this, some studies adopt a compromise by freezing the diffusion model and instead searching for optimized noises (Wagenmaker et al., 2025; Hansen-Estruch et al., 2023), a strategy often referred to as inference-time scaling (Ma et al., 2025b). However, these approaches rely heavily on well-pretrained diffusion policies and are fundamentally limited by their performance upper bound. Another explored direction is to adopt policy gradient methods (such as PPO (Schulman et al., 2017)) for diffusion policy optimization, which models the denoising process as a multi-step Markov decision process (MDP) and uses Gaussian approximations to compute the log-likelihood of intermediate denoising steps (Black et al., 2024b; Ren et al., 2025). However, the crude Gaussian-based approximation only provides reasonable likelihood information when adopting sufficiently small denoising steps, which inevitably results in large exploration spaces and prolonged training, making such methods difficult to scale and prone to approximation error accumulation in practice. Therefore, a critical research question arises: Can we build a more effective and stable RL method for diffusion policy optimization?
@@
To answer this question, we turn our attention to the KL-regularized RL objective, which offers a nice, closed-form weighted regression objective for optimal policy extraction (Peng et al., 2019). We can thus optimize a diffusion policy by incorporating an exponential reward- or value-based weighting term, scaled by a temperature parameter, into the standard diffusion regression loss (Lee et al., 2023; Kang et al., 2023; Zheng et al., 2024). Although promising, this approach also suffers from several limitations. A fundamental issue is that weighted regression can only achieve greedy reward maximization when the temperature parameter is set to a large value, which easily leads to exploding loss and training instability. Moreover, the learning loss becomes dominated by a small number of high-reward samples, which severely undermines training effectiveness and scalability even with increased data (Park et al., 2024). To address the previous challenges, we propose DIPOLE (Dichotomous diffusion Policy improvement), a novel RL framework designed for highly stable and controllable diffusion policy optimization. Specifically, we introduce a greedified KL-regularized RL objective, which regularizes policy learning towards a value-reweighted reference policy. Interestingly, we show that the original unstable exponential weighting term in the optimal policy can be decomposed into two bounded smooth dichotomous terms. This naturally allows us to decompose the optimal policy into a pair of stably learned dichotomous policies: one aims at reward maximization and the other focuses on reward minimization. Moreover, the optimized policy can be recovered through a linear combination of the scores from both dichotomous policies, which closely aligns with the widely used classifier-free guidance mechanism in diffusion models (Ho and Salimans, 2022), enabling perfect controllability over the greediness of action generation.
@@
Extensive experimental results demonstrate the effectiveness of DIPOLE across a wide range of locomotion and manipulation tasks in ExORL (Yarats et al., 2022) and OGBench (Park et al., 2025a) benchmarks, evaluated under both offline and offline-to-online RL settings. Furthermore, we scale our learning approach to a large vision-language-action (VLA) model and evaluate it on the large-scale real-world autonomous driving benchmark NAVSIM (Dauner et al., 2024), showcasing significant performance improvements over the pre-trained baseline. These results highlight the strong applicability of DIPOLE for complex, real-world decision-making scenarios. Project Page: https://lrmbbj.github.io/DIPOLE/
@@ 2 Preliminary
Diffusion/flow matching policies. Diffusion and flow matching models have attracted significant attention due to their strong expressiveness in capturing multi-modal data distributions, making them popular policy classes for complex decision-making tasks such as robotics (Chi et al., 2023; Black et al., 2024a) and autonomous driving (Zheng et al., 2025c; Tan et al., 2025). The action generation can be formulated as a state-conditional generation problem in which a probability path transforms a source distribution (typically a standard Gaussian) into a target action distribution. A neural network $\epsilon_{\theta}$ is trained to predict the noise along the path using the objective over a given dataset $\mathcal{D}$:
@@
where $a_{t}=\alpha_{t}a+\sigma_{t}\epsilon$ (we use the subscript $t$ to distinguish diffusion steps from MDP steps $k$), with $\alpha_{t}$ and $\sigma_{t}$ being predefined noise schedules commonly used in score-based diffusion models (Song et al., 2021) or flow matching models (Lipman et al., 2022). The multi-step diffusion process endows diffusion models with strong distribution-fitting capabilities. However, it also poses challenges for RL fine-tuning: gradient propagation through the entire diffusion process is costly and unstable; the exact likelihood computation with diffusion models is intractable, causing a series of problems when optimizing with existing policy gradient RL methods due to approximation error.
@@ 3.1 KL-Regularized Objective in RL
where $\beta>0$ is the temperature parameter, and $D_{\text{KL}}\left(p\|q\right)=\mathbb{E}_{x\sim p}\left[\log\left(p(x)/q(x)\right)\right]$. $G(\cdot)$ is the evaluated return to be maximized, which can either be the reward function $r(s,a)$ as in single-step problems such as LLM RL fine-tuning (Korbak et al., 2022; Shao et al., 2024), or the action-value function $Q^{\pi}(s,a)$ or advantage function $A^{\pi}(s,a)$ as in standard multi-step settings. The specific choice of reference policy $\mu$ gives rise to different RL task settings. For example, setting $\mu$ to be the uniform distribution, we recover maximum entropy RL as in SAC (Haarnoja et al., 2018); setting $\mu$ to be the behavior policy in offline datasets $\mathcal{D}$, we obtain many offline RL algorithms (Wu et al., 2019; Xu et al., 2023a; Garg et al., 2023); lastly, setting $\mu$ to be a pre-trained policy $\pi_{0}$ or the recently updated policy $\pi_{k-1}$, it corresponds to offline-to-online fine-tuning scenarios (Nakamoto et al., 2023; Li et al., 2023) or trust-region style online policy optimization (Schulman et al., 2015).
The flexibility of the KL-regularized RL framework makes it an ideal choice for diffusion policy optimization. The best part is, it is known that the optimization objective in Eq. (2) also provides a closed-form solution for optimal policy $\pi^{\star}$ as follows (Nair et al., 2020):
@@
Intuitively, the optimal policy is a reweighted version of the reference policy $\mu$, in which actions with higher values are assigned greater probability density. As shown in several existing studies  (Kang et al., 2023; Zheng et al., 2024), if given a pre-trained diffusion policy $\epsilon_{\theta}$ trained with Eq. (1) as the reference policy $\mu$, we can further optimize it with the weighted diffusion loss in Lemma 1 to extract the optimal diffusion policy $\epsilon^{\star}$.
@@
We can generate optimal $a\sim\pi^{\star}(a|s)$ in Eq. (3) by optimizing the weighted diffusion loss in Eq. (4) and solving the diffusion reverse process with obtained $\epsilon^{\star}$ (Zheng et al., 2024).
@@
Compared to diffusion-based RL methods that rely on unstable reward/value maximization (Xu et al., 2023b; Clark et al., 2023) or biased likelihood approximation (Black et al., 2024b; Ren et al., 2025), Eq. (4) offers a simple and scalable training scheme for policy optimization, requiring only the addition of a weighted term to the base diffusion learning objective in Eq. (1). Despite its simplicity, we do not observe the adoption of this scheme in many recent diffusion-based RL methods. Why is that? Actually, there exist several limitations for this exp-weighted regression scheme:
  • Optimality-stability trade-off. As the exponential function $\exp(\cdot)$ grows rapidly, a high-quality action with a large $G(s,a)$ value can lead to an extremely large weight term when $\beta$ is large, causing the explosion of learning loss and destabilizing the training process (illustrated in Figure 1). In practice, many methods mitigate this issue by either using a small $\beta$ or clipping the weighting term (Garg et al., 2023; Xu et al., 2023a; Hansen-Estruch et al., 2023). However, these treatments compromise the optimality of the extracted policy.
  • Inefficient learning. The training loss becomes dominated by a small number of high-return samples, which is inefficient for policy optimization (Park et al., 2024). Additionally, poor-quality samples still retain positive weight, which can adversely affect policy learning. The constrained optimization objective also makes the learning process highly dependent on the quality of the reference policy $\mu$, thereby limiting the potential for greedy policy optimization.
@@ 3.2 Dichotomous Diffusion Policy Improvement
In this revised objective, we instead regularize policy $\pi$ with a greedified, value-aware reference policy weighted by $\color[rgb]{0.76953125,0.1953125,0.19921875}\sigma\left(\beta G(s,a)\right)/Z(s)$}, where $Z(s)$ denotes the normalization factor and $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. This design shares a similar spirit with some offline RL methods that enhance policy performance by regularizing towards a greedier behavior policy or reward-weighted datasets (Singh et al., 2022; Hong et al., 2023; Xu et al., 2025). It is worth noting that we use a bounded and smooth sigmoid function as the weighting function, which greedily assigns high weights to high-return samples while avoiding numerical instability. Moreover, we introduce a new hyperparameter $\color[rgb]{0.76953125,0.1953125,0.19921875}\omega$}, termed the greediness factor, which provides an additional interface for adjusting the greediness of policy extraction. We will reveal its role in the later derivation. Based on the optimization objective in Eq. (5), we can get its closed-form solution as follows:
@@
where $C$ is a constant. This shows that the score function of the optimal policy $\pi^{\star}$ can be expressed as a linear combination of scores of the dichotomous policies, weighted by $\omega$. Due to the inherent connection between the score function and the noise predictor in diffusion model (Ho et al., 2020), we can use $\tilde{\epsilon}\left(a_{t},s,t\right)=(1+w)\epsilon^{+}_{\theta_{1}}\left(a_{t},s,t\right)-w\epsilon_{\theta_{2}}^{-}\left(a_{t},s,t\right)$ in the reverse process of diffusion or flow matching for action sampling.
Interestingly, the formulation in Eq. (3.2) is remarkably similar to classifier-free guidance (CFG) (Ho and Salimans, 2022), a popular method for enhanced conditional diffusion generation, which has the form of $\tilde{\epsilon}(x_{t},c,t)=(1+\omega)\epsilon_{\theta}(x_{t},c,t)-\omega\epsilon_{\theta}(x_{t},t)$, where $\epsilon_{\theta}(x_{t},c,t)$ is a conditioned version of $\epsilon_{\theta}(x_{t},t)$ with conditioning signal $c$. This reveals the inherent connection between our greedified KL-regularized RL objective and the CFG mechanism. Intuitively, our method further strengthens the positive distribution by pushing the negative distribution in the opposite direction, thus enabling flexible control of the optimality level of generated actions with the greediness factor $\color[rgb]{0.76953125,0.1953125,0.19921875}\omega$} (see illustration in Figure 1). Our final formulation also has some similarity with CFGRL (Frans et al., 2025), which can be perceived as setting $\pi^{+}\propto\mu\cdot\mathbb{I}_{A\geq 0}$ and $\pi^{-}=\mu$ (AA is the advantage function). However, their method lacks theoretical backing, and using identical weights for both positive and negative samples limits the greediness of policy optimization, leading to suboptimal performance.
@@ 3.3 Practical Implementations
End-to-end autonomous driving. We also implement DIPOLE to train a large end-to-end autonomous driving model to demonstrate its scalability to solve real-world complex tasks. Specifically, we employ a non-reactive pseudo-closed-loop simulation based on real-world datasets for policy training. In this setup, the return $G(s,a)$ is defined by a reward function that evaluates trajectory quality based on safety, progress, and comfort. We employ a vision-language model (Florence-2 (Xiao et al., 2024)) as the encoder and a diffusion action head as the decoder (Zheng et al., 2025c). The model processes images from the left-front, front, and right-front cameras, along with language instructions such as ”turn left”, ”turn right”, and ”go straight”. This architecture results in a 1-billion parameter model, which we name DP-VLA, and is pre-trained using imitation learning. Subsequently, two separate LoRA modules are applied to the decoder to construct the positive and negative policies, allowing us to leverage Eq. (9) for training. We follow the offline-to-online RL setting to fine-tune the VLA model. Further implementation details are provided in Appendix E.